OpenStudy (anonymous):

4 year later my age will be twice of my cousin.4 years before i was 6 times his age. what is my age.

OpenStudy (anonymous):

a: your age b: your cousin's age 4 years later, my age will be twice that of my cousin's: (1) a+4=2(b+4) 4 years before, I was 6 times his age: (2) a-4=6(b-4) From (1): a+4=2b+8 (3) a=2b+4 Plugging in (3) into (2): (2b+4)-4=6(b-4) 2b=6b-24 -4b=-24 (4) b=6 Plugging in (4) into (3): a=2(6)+4 (5) a=16 From (4) and (5): Your age is 16 and your cousin's age is 6.